OpenStudy (adamaero):

A wavefunction for a particle confined in some region of space is given by the expression: φ(x)=Asin^2(αx)+Bx^(-2) What are the dimensions (i.e., units) of A and B?

OpenStudy (irishboy123):

@osprey

OpenStudy (adamaero):

Sin(x) is odd: f(–x) = –f(x) A squared function is even. Wolfram Alpha says it has even parity...whatever that means. https://www.wolframalpha.com/input/?i=y+%3D+sin%5E2(x)%2BBx%5E-2 The graph looks like this: https://www.wolframalpha.com/input/?i=plot+sin%5E2(x)%2Bx%5E-2

OpenStudy (killemall):

Wave functions are used to show probabilities, so they don't have units.

OpenStudy (irishboy123):

\(\large \Psi (x)=Asin^2(\alpha x)+\dfrac{B}{x^2}\), right?! and you're gonna normalise this as \(\large \int\limits_{-\infty}^{\infty} |\Psi|^2 dx = 1\) ....which is unitless using dimensional analysis... the integration amounts to \(\Sigma |\Psi|_i^2 ~ dx_i\), and being unitless means that \(\Psi\) has units \(\dfrac{1}{\sqrt{m}}\) as \(dx_i\) has units \(m\) A must have the same units as sine is unitless \(\dfrac{B}{m^2} \equiv \dfrac{1}{\sqrt{m}}\) means B has units \(m^{3/2}\) so i reckon ...